how many moles are there in 2.30 x10^24 atoms of silver



1
Expert's answer
2021-01-11T03:49:05-0500

Solution:

The Avogadro constant NA= 6.02×1023 units in one mole of any substance. The units may be electrons, atoms, ions, or molecules.

The number of particles of a substance can be calculated using the Avogadro constant NA, the amount of substance in mole n.

Number of particles N = Avogadro constant NA × amount n

Therefore,

n = 2.30×1024 / 6.02×1023 = 3.82 mol

Answer: there are 3.82 moles of Ag.